# Client Clustering Meets Knowledge Sharing: Enhancing Privacy and Robustness in Personalized Peer-to-Peer Learning

### v1

@@ V-C Tolerating Poisoning Attacks
Figure 4: Impact of attacks on P4 without secure aggregation before grouping and after grouping for three non-IID datasets and a linear model under 30% of malicious clients.
@@
Figure 6 presents the results averaged across multiple seeds. Following the used benchmark [73], an accuracy drop below 10% indicates good defense performance. Based on this threshold and results across both datasets, P4 with the proposed secure aggregation tolerates up to 50% malicious clients under label flipping, 30% under byzantine-zero, 40% under byzantine-random, and 30% under byzantine-flip. The likely reason for the high attack impact above 30% of malicious clients is that m-Krum performs well only when the number of malicious clients is below $(n-2)/2$, according to [36].
@@ D-B Robustness
(a) Varying # of clients for grp. size = 8
@@ F-B Assumptions
[…]

### v2

@@ V-C Tolerating Poisoning Attacks
Figure 4: Attack impact on P4 without secure aggregation before- and after-grouping for 3 non-IID datasets and a linear model with 30% malicious clients.
@@
Figure 6 presents the results averaged across multiple seeds. Following the used benchmark [73], an accuracy drop below 10% indicates good defense performance. Based on this threshold and results across both datasets, P4 with the proposed secure aggregation tolerates up to 50% malicious clients under label flipping, 30% under byzantine-zero, 40% under byzantine-random, and 30% under byzantine-flip. The likely reason for the high attack impact above 30% of malicious clients is that m-Krum performs well only when the number of malicious clients is below (n2)/2(n-2)/2, according to [36].
@@ D-B Robustness
(a) Varying the number of clients for group size = 8
@@ F-B Assumptions
    $\mathbb{E}\bigl[\|\epsilon_{\mathrm{clip}}\|_{1}\bigr]\leq\mathbb{E}\Bigl[\|g\|_{1}\,\mathbf{1}_{\{\|g\|_{2}>\mathcal{C}\}}\Bigr],$